# On automatic boundedness of operators from ordered to topological vector spaces

### 16 Apr 2026 v3

@@ Abstract
We study continuity and topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces. The Levi operators and Lebesgue operators are investigated.
@@ 1 Introduction and preliminaries
Automatic boundedness of operators enjoying some additional properties plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g. [7, 8]). Another case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 6]. Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators. The present paper is devoted to the question how far one can go with other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
Throughout, vector spaces are real, and operators are linear. We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively). Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS. It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous. In [15, Theorem 2.1], this result is extended to the OBS setting.
Let $T:X\to Y$ be an operator from an NS $X$ to a TVS $(Y,\tau)$. In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$. For example, using of finite sets gives no proper subclass, because $T$ always carries finite sets into $\tau$-bounded sets. On the other hand, usage of compactness returns us to bounded operators. Indeed, if $T$ is $\tau$-bounded then $T$ takes compact sets into bounded sets. On the other hand, if $T$ is not $\tau$-bounded then there exist a circled $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ such that $Tx_{n}\notin n^{2}U$ for all $n\in\mathbb{N}$. In this case, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ under $T$ is not $\tau$-bounded.
@@
Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it is bounded along normalized (unconditional) basic sequences of $X$?
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded along bounded unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space). Therefore, it makes sense to deal with the following questions.
Question 1.2.
Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
Question 1.3.
How to describe the class of Banach lattices such that every linear functional on a Banach lattice bounded along bounded disjoint sequences is bounded?
An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to this class.
The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain. We mainly focus on Banach spaces in the domain and (when suitable) on Fréchet spaces. The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
This paper is organized as follows. In Section 2, we study order-to-topology continuous/bounded operators between OBSs (or, also, OVSs) and TVSs. Conditions providing continuity of such operators are investigated. Section 3 is devoted to conditions providing boundedness/continuity of the Levi operators and the Lebesgue operators. For instance, Theorem 3.3 tells us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded. Theorem 3.12 states that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded.
For further unexplained terminology about TVSs and OVSs, we refer the reader to [1, 3, 4].
2 On boundedness of some operators between ordered TVSs
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]). Some of their results were extended to operators from an OBS to a TVS in [12]. It is important to find further conditions on operators in ordered TVSs which provide automatic continuity/boundedness.
In this section, we continue the research initiated in [12, 13, 11, 15]. Propositions 2.3, 2.4, and 2.5, provides several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
Then, we study $\tau$-Levi and $\tau$-Lebesgue operators following [5, 19, 12] in more general OVS-setting in the domain. Lemma 3.11 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone. Theorem 3.12 states that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded. Theorem 3.4 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
We need several further notions and notation. Recall that a net $(x_{\alpha})$ in an OVS $X$
@@ 2 On boundedness of some operators between ordered TVSs
    order converges to $x$ (o-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
@@
    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
  3. -
    is order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
  4. -
    is eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
  5. -
    is o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})\operatorname{\xrightarrow[]{ru}}0$).
It is straightforward to see that every o-convergent and every ru-convergent (hence, o-Cauchy and ru-Cauchy) net is eventually order bounded. In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
@@ 1. $a)$
    order bounded if $T[a,b]$ is order bounded for every $a,b]\subseteq X$ ((shortly, $T\in{\cal L}_{ob}(X,Y))$.
@@ 2. $b)$
    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly, $T\in{\cal L}_{oc}(X,Y))$.
@@ 3. $c)$
    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ ((shortly, $T\in{\cal L}_{rc}(X,Y))$.
@@ 1. $d)$
    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{onc}(X,Y)$, when $Y$ is an NS)).
@@ 2. $e)$
    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write $T\in{\cal L}_{rnc}(X,Y)$, when $Y$ is an NS)).
@@ 3. $f)$
    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write $T\in{\cal L}_{onb}(X,Y)$, when $Y$ is an NS)).
@@ 1. $g)$
    continuous $(T\in{\cal L}_{c}(X,Y))$ if $Tx_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
@@ 2. $e)$
    topologically-bounded $(T\in{\cal L}_{tb}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$ ((we write ${\cal L}_{b}(X,Y)$, when $Y$ and $Y$ are NSs)).
Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{onc}(X,Y)$, etc. given by Definition 2.1 are vector spaces.
We recall several notions on collectively qualified sets of operators (cf. [9, 10, 12, 13, 17, 18, 15]). A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}\operatorname{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
@@ 1. $b)$
    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
@@ 2. $c)$
    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$, when $Y$ is an NS)).
@@ 3. $d)$
    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$, when $Y$ is an NS)).
@@ 1. $e)$
    equi-continuous $({\cal T}\in\text{\bf L}_{c}(X,Y))$ if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
@@ 2. $f)$
    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{tb}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded ((we write ${\cal T}\in\text{\bf L}_{b}(X,Y)$, when $X$ and $Y$ are NSs)).
The sequential versions of Definition 2.1 b)–e), g) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
In many cases, equi-order-to-topology bounded sets agree with equi-ru-to-topology continuous ones.
@@
([12, Theorem 2.1]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS. If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$. The following proposition is a topological counterpart of [15, Theorem 2.1].
@@
([12, Theorem 2.3]) $\text{\bf L}_{otc}(X,Y)\subseteq\text{\bf L}_{otb}(X,Y)$, whenever $X$ is an Archimedean OVS with a generating cone, and $Y$ is a TVS.
In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded. The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
Proposition 2.5.
([12, Theorem 2.5]) $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [12], we also collect few more statements below.

### 20 Apr 2026 v4

# On automatic continuity of operators from ordered to topological vector spaces

@@ Abstract
We study continuity and boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered to topological vector spaces. Several results on automatic continuity of operators from ordered Fréchet spaces to topological vector spaces are included. Levi and Lebesgue operators especially are investigated.
@@ 1 Introduction and preliminaries
Automatic continuity of operators enjoying some additional properties plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g. [7, 8]). Another case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 6]. Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators. This paper is devoted to the study of the question how far one can go with other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
Throughout, vector spaces are real, and operators are linear. When suitable, we abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively). In what follows, symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval in an OVS $X$; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS. It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous. In [15, Theorem 2.1], this result is extended to the OBS setting. In [12], the further extension is obtained in the case of a TVS in the co-domain. In the present paper, we prove Theorem 2.6 that generalizes [12, Theorem 2.1] to operators from an ordered Fréchet space with a closed generating cone to a TVS.
Let $T:X\to Y$ be an operator from a normed space $X$ to a TVS $(Y,\tau)$. In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$. For example, using of finite sets gives no proper subclass, because $T$ always carries finite sets into $\tau$-bounded sets. On the other hand, usage of compactness returns us to bounded operators. Indeed, if $T$ is $\tau$-bounded then $T$ takes compact sets into bounded sets. On the other hand, if $T$ is not $\tau$-bounded then there exist a circled $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ such that $Tx_{n}\notin n^{2}U$ for all $n\in\mathbb{N}$. In this case, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ under $T$ is not $\tau$-bounded.
@@
Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it takes normalized (unconditional) basic sequences of $X$ onto bounded sequences of $Y$?
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded on normalized unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space). Therefore, it makes sense to deal with the following questions.
@@
    Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ that is bounded on normalized unconditional basic sequences of $X$ is bounded?
@@
    How to describe the class $S$ of Banach lattices such that every linear functional on a Banach lattice $X$ from $S$ is bounded, whenever it is bounded on every normalized disjoint sequence of $X$? An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to the class $S$.
The present paper deals with operators between vector spaces, where the order structure is assumed in the domain or/and codomain. We focus on Banach spaces and, when suitable, Fréchet spaces in the domain. The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
This paper is organized as follows. In Section 2, we survey briefly some recently published results on order-to-topology continuous/bounded operators between OBSs and TVSs. Then, we prove Theorem 2.6 on continuity of order-to-topology bounded operators from an ordered Fréchet space with a closed generating cone to a TVS. In Section 3, we study conditions providing continuity of Levi and Lebesgue operators. For instance, we prove: Theorem 3.2 telling us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is continuous; and Theorem 3.12 saying that each $\sigma$-w-Lebesgue operator from an ordered Banach space with a closed generating normal cone to a normed space is bounded.
For further unexplained terminology and notation, we refer the reader to [1, 3, 4].
2 On continuity of some operators between ordered topological vector spaces
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]). Some of their results were extended to operators from an OBS to a TVS in [12]. It is important to find further conditions on operators in ordered TVSs which provide automatic continuity/boundedness. In this section, we continue the research initiated in [12, 13, 11, 15]. First, we point out Propositions 2.4 and 2.5 providing several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators. Then, we prove Theorem 2.6 telling us that every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is continuous. We need several further notions and notation.
Definition 2.1.
A net $(x_{\alpha})$ in an ordered vector space $X$ is
  1. $a)$
    order convergent to $x$ (o-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
  2. $b)$
    relative uniform convergent to $x$ (ru-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
  3. $c)$
    order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
  4. $d)$
    eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
  5. $e)$
    o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{ru}}0$).
It is straightforward to see that every o-convergent and every ru-convergent (hence, every o-Cauchy and every ru-Cauchy) net is eventually order bounded. In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
@@ 1. $a)$
    order bounded $(T\in{\cal L}_{ob}(X,Y))$ if $T[a,b]$ is order bounded for every $a,b]\subseteq X$.
@@ 2. $b)$
    order continuous $(T\in{\cal L}_{oc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{o}}0$.
@@ 3. $c)$
    ru-continuous $(T\in{\cal L}_{rc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{ru}}0$.
@@ 1. $d)$
    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
@@ 2. $e)$
    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
@@ 3. $f)$
    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$.
@@ 1. $g)$
    continuous $(T\in{\cal L}_{tc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
@@ 2. $e)$
    topologically-bounded $(T\in{\cal L}_{b}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$.
Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{otc}(X,Y)$, etc. given by Definition 2.2 are vector spaces.
We recall several notions on collectively qualified sets of operators (cf. [9, 10, 12, 13, 17, 18, 15]). A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
@@ 1. $b)$
    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
@@ 2. $c)$
    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
@@ 3. $d)$
    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$.
@@ 1. $e)$
    equi-continuous $({\cal T}\in\text{\bf L}_{tc}(X,Y))$ if ${\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
@@ 2. $f)$
    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{b}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded.
The sequential versions of Definition 2.2 b)–e), g) and Definition 2.3 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{tc}^{\sigma}(X,Y)$, etc.).
We recall several recent results concerned the notions listed in Definitions 2.2 and 2.3.
@@
([12, Theorem 2.1]) If $X$ is an OVS and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$. If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
In particular, each order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, then ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$. The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
@@
([12, Theorem 2.5]) If $X$ is an OBS with a closed generating cone and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$.
It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [12], we collect few more statements below.